The average number of milligrams of cholesterol in a cup of a certain


brand ice cream is 670 mg and standard deviation is 45 mg, Assume the


variable is normally distributed.


a. If a cup of ice cream is selected, what is the probability that the


cholesterol content will be more than 700 mg?


b. If a sample of 20 cups is selected what is the probability that the mean of


the sample will be larger than 700mg?

1
Expert's answer
2022-03-29T17:47:14-0400

"\mu=670,\sigma=45"

a.

"P(X>700)=P(Z>\frac{700-670}{45})"

"=P(Z>0.67)"

"=1-P(Z<0.67)"

"=1-0.7486"

"=0.2514"


b.

"n=20"

"P(X>700)=P(Z>\frac{700-670}{\frac{45}{\sqrt{20}}})"

"=P(Z>2.98)"

"=1-P(Z<2.98)"

"=1-0.9986"

"=0.0014"